in a certain triangle the measure of one angle is double the measure of a second angle but is 10 degrees less than the measure of the third angle. [the sum of the measures of the three interior angles of a triangle is always 180 degrees.] what is the measure of each angle?: in a certain triangle the measure of one angle is double the measure of a second angle but is 10 degrees less than the measure of the third angle. [the sum of the measures of the three interior angles of a triangle is always 180 degrees.] what is the measure of each angle?
Answer by vleith(1174)

Let X be the first angle. You are told the second angle is X/2. The third angle is X+10.
X + X/2 + (X+10) = 180
5X/2 = 170
5X = 340
X = 68
That means the second angle is 68/2 = 34
The third angle is 68+10 = 78
Do those three total to 180?